The sum of four consecutive whole numbers are 170. Find the largest number
Answer by jim_thompson5910(35256)

The sum of four consecutive whole numbers are 170. Find the largest number

--------------------------------------------------------------------------------------------------------------

x + (x+1) + (x+2) + (x+3) = 170

x + x+1 + x+2 + x+3 = 170

4x + 6 = 170

4x = 170-6

4x = 164

x = 164/4

x = 41

----------------------

x = 41

x+1 = 41+1 = 42

x+2 = 41+2 = 43

x+3 = 41+3 = 44

----------------------

The four numbers are 41,42,43,44. The largest is 44